OpenStudy (anonymous):

Determine the convergence or divergence. If the sequence converges, find the limit.

OpenStudy (anonymous):

\[\left\{ an \right\} = \left\{ 1+\frac{ (-1)^{n} }{ n } \right\}\]

OpenStudy (anonymous):

not sure how to go about starting this. would i let {an}=f(x) then try to solve the limit?

OpenStudy (experimentx):

the sequence converges, why?? a0<a2< ... a3<a1

OpenStudy (anonymous):

i know if it was just(-1)^n/n it would converge by the absolute theorm but, im not sure what the additional 1 would do.

OpenStudy (anonymous):

because f(x) is not defined on [1, infinity)

OpenStudy (anonymous):

...???

OpenStudy (experimentx):

sorry ... i was busy :((

OpenStudy (experimentx):

notice that is is squeezed between odd terms and even terms

OpenStudy (anonymous):

so its the squeeze theorm?

OpenStudy (anonymous):

why is it squeezed?

OpenStudy (experimentx):

somewhat like that \[ a_{2n}>a_{2n-2}> ..... >a_{2n-1}>a_{2n+1}\]

OpenStudy (experimentx):

sorry ,, i put in opposite order \[ a_{2n+2}>a_{2n}> ..... >a_{2n+1}>a_{2n-1} \]

OpenStudy (experimentx):

again you need to concept of limit point. do you know monotone convergence theorem?

OpenStudy (anonymous):

we haven't done that. we've done squeeze, absolute, test for convergence, nth term test for divergence, convergence of a geometric sequence, limits of functions, one with l'hopital rule.

OpenStudy (experimentx):

just show that \[ \lim_{n\to\infty } |1 + (-1)^n/n - 1| = 0\] it is enough to prove the limit of sequence is 1.

OpenStudy (anonymous):

so it is the absolute value theorem'

OpenStudy (experimentx):

no ... definition of limit

OpenStudy (anonymous):

where are you getting "n-1" from?

OpenStudy (experimentx):

haa ... looks like you didn't get it